### v1

@@ 6.5.2 Microlensing event detection
The microlensing method is currently the best method for detecting “cold” exoplanets and rogue planets (or free floating planets, FFPs) that are not bound by the gravity of stars [131, 132]. Since 2003, the microlensing method has discovered more than 200 exoplanets. These planets occupy relatively unique positions in parameter space, greatly improving our understanding of “cold” exoplanets and rogue planets [133, 134].
@@
Based on the known luminosity distribution function of bulge stars and the currently estimated rogue planet mass distribution function [135], we expect that a CSST total observation time (i.e. exposure plus readout time) of 15 days will be able to detect about 60 microlensing events caused by rogue planets of Mars mass (see Figure 12), and dozens of microlensing events caused by rogue planets of larger mass.
We also expect to detect about 50 signals from “cold” planets gravitationally bound to their host stars. These “cold” planets are about $1-30$ AU from their host stars and have masses as low as about Mars mass. This type of exoplanet is almost impossible to detect by other detection methods or ground-based microlensing surveys, and they can play an important role in testing the core accretion model of planet formation theory. It is expected that these extremely low-mass rogue planets and “cold” planets that CSST can detect will greatly expand the known planetary parameter space, and have very important scientific value for understanding the formation and evolution of planetary systems.
@@ 6.5.3 Transiting exoplanet system detection
As the main method for discovering exoplanets, the transit method has the characteristic of batch searching for exoplanets with high efficiency. At present, nearly 5,000 exoplanets have been confirmed by the transit method, providing a rich sample for statistical research on exoplanets and revealing various dynamic and physical characteristics of exoplanet systems. All known transiting planets are mainly in the Kepler field, as well as in the solar neighborhood. The basic principle of the transit method is to measure the tiny periodic photometric changes caused by the planet passing over the surface of the star through high-precision time-series photometric observations, thereby obtaining orbital characteristics, such as planet radius and orbital period, which helps us understand the properties of exoplanets.
The transit method for searching exoplanets requires long-term observation of a large number of stars in the same area of the sky. CSST has a large FoV and deep magnitude limit, which can monitor many stars simultaneously and search for transit signals. Meanwhile, CSST has a high angular resolution and can observe areas with dense star fields, such as the vicinity of the galactic bulge and star clusters. With a larger aperture than Kepler, CSST can detect transiting planets around farther and fainter bulge stars, which can reveal the planet formation mechanism in different environments throughout our galaxy. We find that with an exposure of 300 seconds, the CSST can measure stars with a photometric accuracy of 0.5% for magnitudes below 20, and its detection depth is higher than that of current space telescopes which use the transit method to detect planets, such as the Kepler apace telescope and TESS. With an observational time of 7 days, more than 100 transiting exopalnet candidates can be detected.
@@ 6.5.4 Exoplanet atmosphere
Transit spectroscopy, direct imaging and Doppler spectroscopy are the best methods for studying the atmospheres of exoplanets. Although planets are too dark and too small, their size is not that different from their host stars. When a planet moves in front of a star, the star light is partially blocked, so it appears to be darker, and the degree of darkening depends on the wavelength. The “transit method” is continuously tracking stars, searching for the phenomenon of periodic dimming of stars and measuring the degree of dimming of planetary systems in multiple bands, which can be used to detect planets and obtain the atmospheric spectrum of planets. When a planet moves behind its host star, the radiation of the planet is completely blocked. By comparing the total radiation of the entire system at this time with that before the planet is blocked, the radiation of the planet itself or its reflection can be obtained. This is called the “secondary eclipse method”.
The CSST-SC includes a slitless spectrometer with a resolution of about 200 and a wavelength coverage of $\sim$0.25 to 1 $\mu$m, which can be used to obtain the transmission or emission spectra of planetary atmospheres. The CSST-IFS can provide a resolution of $\geq 1000$, a wavelength coverage of $0.35-1.0$ $\mu$m transmission or reflection spectra, and can directly distinguish some wide-spaced planets. The CSST-MCI can provide three-color simultaneous high-precision light curves. These instruments have their own advantages and can be used for observation and research of exoplanet atmospheres.
Considering the CSST instruments and band coverage, cooperated with other instruments (e.g. JWST, HST, etc.), we expected to obtain transit spectra and multi-color phase curves of a total of $20-30$ hot and ultra-hot Jupiters, and perform detailed studies of their atmospheric properties. We will also conduct atmospheric studies of planets around K-type and M-type dwarfs, and use CSST to conduct follow-up optical spectroscopy or multi-color observations of exoplanets that have been observed by HST or JWST near-IR transmission spectroscopy. Besides, we will probe about 10 hydrogen ocean planets, super-Earths and warm planets, and use the IFS to perform direct spectroscopic detections of a wide range of young giant planets. Moreover, the CSST-SC spectrometer and MCI should enable effective detection of the atmospheres of most hot Jupiters, some super-Earths and wide-spaced young planets.
@@ 6.6.3 Physical properties and activity of comets
Comets have a wide variety of morphological features during their orbits. CSST has a large FoV and high resolution, making it very suitable for studying the morphology and activity of comets. We also can use the CSST main survey to carry out observations of comets, and build a database of comet physical parameters by combining other survey data. Besides, The short-time scale ($5-30$ s) exposure full-band observation/monitoring mode of the CSST-MCI can measure the light variation of comets with high spatial resolution at the minute level, which can be used to study the rotation period and activity evolution of comets. The start and end positions of comet activity are one of the important parameters for comparing comet activity. Since the imaging depth of the XDF observation of CSST-MCI can reach 30 magnitude in optical band, CSST-MCI can greatly expand the heliocentric distance at which comets are active.
@@ 6.6.5 Discovery and physical properties of natural satellites
Most of the natural satellites of the large planets are irregular satellites, which are usually very dim ($>20$ mag) and difficult to observe using ground-based telescopes [143, 144, 145, 146, 147, 148]. The study of these faint satellites is of great significance to the origin and evolution of the Solar System, and also plays a key role in the navigation of deep space exploration.
@@ 6.7.2 Maintaining and updating the $Gaia$ catalog
The $Gaia$ mission delivers star catalog with unprecedented astrometric precision -- approximately 1 milliarcsecond for objects in the 18 to 21 magnitude range [155]. However, the accuracy of these catalogs, particularly in terms of stellar position, deteriorates over time due to the increasing uncertainty in astrometric parameters. Incorporating new observations is therefore essential to maintain, or even enhance, the initial level of precision [156].
The CSST observations cover $Gaia$ targets of $18-21$ mag, and the combination of common target astrometric observational data between $Gaia$ and CSST can improve the accuracy of the proper motion and parallax solution results of the targets. We will optimally combine $Gaia$ astrometric epoch data with CSST astrometric epoch data, and solve the relevant astrometric parameters in the same set of equations.
@@ 6.8.2 Gamma-ray bursts
Gamma-ray bursts can generally be divided into two categories: long GRBs (duration time scale greater than 2 seconds) and short GRBs (duration time scale less than 2 seconds). The former originates from the core collapse of massive stars, while the latter is produced by the merger of binary compact stars. As the most violent stellar explosion in the Universe, the instantaneous radiation of GRBs is extremely bright. At the same time, gamma-rays have extremely strong penetrability in the ISM. These factors make GRBs detectable even at extremely long distances.
@@ 6.8.3 Fast radio bursts
Fast radio bursts are radio bursts from the distant Universe that last only a few milliseconds but can release as much energy as the Sun releases in a whole day [162]. To date, more than 800 cases have been officially reported, of which about 60 have shown repetitive bursts [163]. In more than 40 cases, the host galaxies and redshifts are determined with arcsecond precision, confirming the cosmological origin of FRBs [164]. However, their astrophysical origin remains a mystery and will remain an important topic in frontier research in the next 5-10 years [165].
@@ 6.8.5 Tidal disruption events
In the CSST era, various ground-based large-field surveys (e.g. the Wide Field Survey Telescope (WFST) and LSST) and X-ray space satellites (e.g. Einstein Probe (EP) X-ray space telescope) have been routinely operated, and the TDE discovery rate will increase by orders of magnitude. CSST has its unique advantages as a space UV and optical telescope, and will discover TDEs in the deep-field survey, follow up important TDEs warned by other surveys, and study the host galaxies of TDEs.

### v3

@@ 6.5.2 Microlensing event detection
The microlensing method is currently the best method for detecting “cold” exoplanets and rogue planets (or free-floating planets, FFPs) that are not bound by the gravity of stars [183, 184]. Since 2003, the microlensing method has discovered more than 200 exoplanets. These planets occupy relatively unique positions in parameter space, greatly improving our understanding of “cold” exoplanets and rogue planets [185, 186].
@@
Based on the known luminosity distribution function of bulge stars and the currently estimated rogue planet mass distribution function [187], we expect that a CSST total observation time (i.e. exposure plus readout time) of 15 days will be able to detect about 20 microlensing events caused by free-floating planets of Mars mass and smaller (see Figure 12), and roughly two dozen more massive free-floating planets.
We also expect to detect about 50 signals from “cold” planets gravitationally bound to their host stars. These “cold” planets are about 1301-30 AU from their host stars and have masses as low as about Mars mass. This type of exoplanet is almost impossible to detect by other detection methods or ground-based microlensing surveys, and they can play an important role in testing the core accretion model of planet formation theory. It is expected that these extremely low-mass rogue planets and “cold” planets that CSST can detect will greatly expand the known planetary parameter space, and have very important scientific value for understanding the formation and evolution of planetary systems.
@@ 6.5.3 Transiting exoplanet system detection
As the main method for discovering exoplanets, the transit method has the characteristic of batch searching for exoplanets with high efficiency. At present, nearly 5,000 exoplanets have been confirmed by the transit method, providing a rich sample for statistical research on exoplanets and revealing various dynamic and physical characteristics of exoplanet systems. All known transiting planets are mainly in the Kepler field, as well as in the solar neighborhood [186]. The basic principle of the transit method is to measure the tiny periodic photometric changes caused by the planet passing over the surface of the star through high-precision time-series photometric observations, thereby obtaining orbital characteristics, such as planet radius and orbital period, which helps us understand the properties of exoplanets.
The transit method for searching exoplanets requires long-term observation of a large number of stars in the same area of the sky. CSST has a large FoV and deep magnitude limit, which can monitor many stars simultaneously and search for transit signals. Meanwhile, CSST has a high angular resolution and can observe areas with dense star fields, such as the vicinity of the galactic bulge and star clusters. With a larger aperture than Kepler, CSST can detect transiting planets around farther and fainter bulge stars, which can reveal the planet formation mechanism in different environments throughout our galaxy. We find that with an exposure of 300 seconds, the CSST can measure stars with a photometric accuracy of 0.5% for magnitudes below 20, and its detection depth is higher than that of current space telescopes which use the transit method to detect planets, such as the Kepler apace telescope and TESS [175, 166]. With an observational time of 7 days, more than 100 transiting exoplanet candidates can be detected.
@@ 6.5.4 Exoplanet atmosphere
Transit spectroscopy, direct imaging and Doppler spectroscopy are the best methods for studying the atmospheres of exoplanets [188]. Although planets are too dark and too small, their size is not that different from their host stars. When a planet moves in front of a star, the star light is partially blocked, so it appears to be darker, and the degree of darkening depends on the wavelength. The “transit method” is continuously tracking stars, searching for the phenomenon of periodic dimming of stars and measuring the degree of dimming of planetary systems in multiple bands, which can be used to detect planets and obtain the atmospheric spectrum of planets. When a planet moves behind its host star, the radiation of the planet is completely blocked. By comparing the total radiation of the entire system at this time with that before the planet is blocked, the radiation of the planet itself or its reflection can be obtained. This is called the “secondary eclipse method”.
[…]
Considering the CSST instruments and band coverage, cooperated with other instruments (e.g. JWST, HST, etc.), we expected to obtain transit spectra and multi-color phase curves of a total of 203020-30 hot and ultra-hot Jupiters, and perform detailed studies of their atmospheric properties. We will also conduct atmospheric studies of planets around K-type and M-type dwarfs, and use CSST to conduct follow-up optical spectroscopy or multi-color observations of exoplanets that have been observed by HST or JWST near-IR transmission spectroscopy. Besides, we will probe about 10 hydrogen ocean planets, super-Earths and warm planets, and use the IFS to perform direct spectroscopic detections of a wide range of young giant planets. Moreover, the CSST-SC spectrometer and MCI should enable effective detection of the atmospheres of most hot Jupiters, some super-Earths and wide-spaced young planets.
@@ 6.6.3 Physical properties and activity of comets
Comets have a wide variety of morphological features during their orbits. CSST has a large FoV and high resolution, making it very suitable for studying the morphology and activity of comets. We also can use the CSST main survey to carry out observations of comets, and build a database of comet physical parameters by combining other survey data. Besides, The short-time scale (5305-30 s) exposure full-band observation/monitoring mode of the CSST-MCI can measure the light variation of comets with high spatial resolution at the minute level, which can be used to study the rotation period and activity evolution of comets. The start and end positions of comet activity are one of the important parameters for comparing comet activity. Since the imaging depth of the XDF observation of CSST-MCI can reach 30 magnitude in optical band, CSST-MCI can greatly expand the heliocentric distance at which comets are active.
@@ 6.6.5 Discovery and physical properties of natural satellites
Most of the natural satellites of the large planets are irregular satellites, which are usually very dim (>20>20 mag) and difficult to observe using ground-based telescopes [196, 197, 198, 199, 200, 201]. The study of these faint satellites is of great significance to the origin and evolution of the Solar System, and also plays a key role in the navigation of deep space exploration.
@@ 6.7.2 Maintaining and updating the $Gaia$ catalog
The $Gaia$ mission delivers star catalog with unprecedented astrometric precision - approximately 1 milliarcsecond for objects in the 18 to 21 magnitude range [208]. However, the accuracy of these catalogs, particularly in terms of stellar position, deteriorates over time due to the increasing uncertainty in astrometric parameters. Incorporating new observations is therefore essential to maintain, or even enhance, the initial level of precision [209].
The CSST observations cover $Gaia$ targets of 182118-21 mag, and the combination of common target astrometric observational data between $Gaia$ and CSST can improve the accuracy of the proper motion and parallax solution results of the targets. We will optimally combine $Gaia$ astrometric epoch data with CSST astrometric epoch data, and solve the relevant astrometric parameters in the same set of equations.
@@ 6.8.2 Gamma-ray bursts
Gamma-ray bursts can generally be divided into two categories: long GRBs (duration time scale greater than 2 seconds) and short GRBs (duration time scale less than 2 seconds) [215]. The former originates from the core collapse of massive stars, while the latter is produced by the merger of binary compact stars. As the most violent stellar explosion in the Universe, the instantaneous radiation of GRBs is extremely bright. At the same time, gamma-rays have extremely strong penetrability in the ISM. These factors make GRBs detectable even at extremely long distances.
@@ 6.8.3 Fast radio bursts
Fast radio bursts are the most energetic cosmic explosions in radio bands [216], typically releasing an immense amount of energy in milliseconds, which can be the same as what the Sun produces in days or even months. To date, more than 860 cases have been officially reported, of which about 8% have shown repetitive bursts [217, 218] and a few are extremely active with hundreds to tens of thousands of bursts from a single FRB [219, 220]. In more than 40 cases, the host galaxies and redshifts are determined with arcsecond precision, confirming the cosmological origin of FRBs [221]. However, their astrophysical origin remains a mystery and will be an important topic in frontier research in the next 5-10 years [222].
@@ 6.8.5 Tidal disruption events
In the CSST era, various ground-based large-field surveys (e.g. the Wide Field Survey Telescope (WFST) [227] and LSST) and X-ray space satellites (e.g. Einstein Probe (EP) X-ray space telescope [228]) have been routinely operated, and the TDE discovery rate will increase by orders of magnitude. CSST has its unique advantages as a space UV and optical telescope, and will discover TDEs in the deep-field survey, follow up important TDEs warned by other surveys, and study the host galaxies of TDEs.